OpenStudy (anonymous):

Order 34 x 102, 1.2 x 107, 8.11 x 10-3, and 435 from least to greatest.

OpenStudy (mertsj):

Are those supposed to be exponents?

OpenStudy (mertsj):

\[34\times10^2, 1.2\times10^7, 8.11\times10^{-3}, 435\]

OpenStudy (anonymous):

34×10^2,1.2×10^7,8.11×10^−3,435 sorry lol

OpenStudy (mertsj):

\[3400, 12000000, .00811, 435\]\[3400,\]

OpenStudy (mertsj):

Put them in order.

OpenStudy (anonymous):

Thanks A LOT

OpenStudy (callisto):

\[34 \times 10^2, 1.2 \times 10^7, 8.11 \times 10^{-3}, 435\] \[435 = 4.35\times 10^2\]\[34\times 10^2 = 3.4 \times 10^3\]So, now, you get \[3.4 \times 10^3, 1.2 \times 10^7, 8.11 \times 10^{-3}, 4.35 \times 10^2\] Rank according to the power.. smaller power of 10 -> smaller value of the number. Oh.. I'm always to late :|